# ML-PWS: Estimating the Mutual Information Between Experimental Time Series Using Neural Networks

### v1

@@ I Introduction
Discussion—We demonstrated how autoregressive sequence prediction models can be trained on time-series data to learn a generative model, which can be combined with PWS to compute the information rate. By applying ML-PWS to a nonlinear model of information processing, we showed that it provides more accurate mutual information estimates than the Gaussian approximation. While this example serves as a proof of concept, it shows the potential of advanced machine learning techniques to automatically derive stochastic models from experimental data, and to enable the computation of information-theoretic measures for complex, high-dimensional data.
@@
While ML-PWS shares similarities with the DoE estimator [30], ML-PWS has the benefit that $\operatorname{\mathrm{P}}(x_{1:n})$ is obtained via an exact marginalisation of $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})\operatorname{\mathrm{P}}(s_{1:n})$ using PWS, and not by training a second neural network as in DoE; this ensures that the marginal statistics $x_{1:n}$ are consistent with the input distribution $\operatorname{\mathrm{P}}(s_{1:n})$, as determined by the setup of the experiment, and the conditional input distribution $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$ of the generative model. Moreover, ML-PWS also enables the efficient computation of the mutual information $I(\boldsymbol{S}^{\prime},\boldsymbol{X}^{\prime})$ between a input signal $\boldsymbol{S}^{\prime}$ with different statistics and the corresponding output $\boldsymbol{X}^{\prime}$, without needing to re-train the marginal model. This is particularly useful if one is interested in the channel capacity, as determined by the input distribution that maximizes the mutual information or information rate for the system of interest. Indeed, in systems without feedback, $\operatorname{\mathrm{P}}(\boldsymbol{x}|\boldsymbol{s})$ is a property of the system and does not change upon changing the input. The generative model then remains the same, such that ML-PWS can directly recompute the mutual information for different input statistics.
In conclusion, ML-PWS represents a promising advance for estimating the mutual information of high-dimensional data, a notoriously difficult problem. By enabling the accurate computation of the mutual information rate, we anticipate that ML-PWS could contribute to a deeper understanding of complex dynamical systems, with potential applications spanning neuroscience, biology, physics, and machine learning.
Acknowledgements.
@@
Appendix A Autoregressive RNN model
@@
$h_{i}=f_{\theta}(s_{i},h_{i-1})$ (10)
where $h_{i}\in\mathbb{R}^{d}$ for $i\in\{1,\ldots,n\}$ and an activation function $f_{\theta}:\mathbb{R}\times\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. The activation function $f_{\theta}$ could for instance be parameterized by a simple neural network layer
$f_{\theta}(s,h)=\tanh(Ws+Vh+b)$ (11)
with parameters $\theta=(W\in\mathbb{R}^{d\times 1},V\in\mathbb{R}^{d\times d},b\in\mathbb{R}^{d})$ and applying $\tanh$ element wise. Other possible choices for $f_{\theta}$ include LSTM units [38] or GRU units [39] which allow the model to better learn long-term dependencies. In any case, given an initial state $h_{0}$ and the parameters $\theta$, a RNN represents a deterministic map from one sequence $s_{1:n}$ to another sequence $h_{1:n}$.
From the sequence $h_{1:n}$ we can obtain a stochastic representation of the output sequence $x_{1:n}$ given an input sequence. To represent a stochastic sequence $x_{1:n}=(x_{1},\ldots,x_{n})$, we decompose the joint distribution as a product of conditional probabilities
Specifically, we extend the recursive relation above by adding a sampling step to obtain $x_{i}$
$h_{i}$ $=f_{\theta}(s_{i},x_{i-1},h_{i-1})$ (13)
$x_{i}\mid h_{i}$ $\sim\mathcal{N}(\mu_{\phi}(h_{i}),\sigma_{\alpha}(h_{i}))$
such that each $x_{i}$ is a normal-distributed random variable whose mean $\mu_{\phi}(h_{i})$ and standard deviation $\sigma_{\alpha}(h_{i})$ are obtained from the current hidden state $h_{i}$.
The autoregressive model represents a generative model for the distribution $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$. Generating $x_{1:n}$ given $s_{1:n}$ is done sequentially. For $i=1,\ldots,n$, the sampling procedure alternates between computing the parameters of the conditional distribution, $\hat{\mu}_{i}=\mu_{i}(x_{1:i-1},s_{1:i})$, $\hat{\sigma}_{i}=\sigma_{i}(x_{1:i-1},s_{1:i})$, and sampling the next $x_{i}\sim\mathcal{N}(\hat{\mu}_{i},\hat{\sigma}_{i})$. Thus, the conditional probability of the resulting output sequence is given by
$\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$ $=\prod^{n}_{i=1}\operatorname{\mathrm{P}}(x_{i}|x_{1:n-1},s_{1:n})$ (14)
@@
In practice we use the following form form for $\mu_{\phi}$ and $\sigma_{\alpha}$
$\mu_{\phi}(h)$ $=W_{\mu}h+b_{\mu}$ (15)
$\sigma_{\alpha}(h)$ $=\exp(W_{\sigma}h+b_{\sigma})$ (16)
where the weights $W_{\mu}$ and $W_{\sigma}$ are $d$-dimensional row vectors and the exponential function ensures that the standard deviation is always positive. For brevity of notation we use $\boldsymbol{\theta}=(\theta,W_{\mu},b_{\mu},W_{\sigma},b_{\sigma})$ to refer to all weights of the model.
The recursive relation (13) defines the conditional probability distribution $P(x_{1:n}|s_{1:n},\boldsymbol{\theta})$ of the output sequence given the input sequence for our generative model. Note that since $h_{i}$ depends on $s_{i}$ and $x_{i-1}$, as well as $h_{i-1}$, it encodes information about the entire past $s_{1:i}$ and $x_{1:i-1}$ and thus the conditional probability of the sequence $x_{1:n}$ is given by Eq. 14.
A.1 Training
There are a few practical considerations for efficiently training the model. Training is performed in iterations and it is often beneficial to introduce stochasticity between iterations to speed up gradient descent and regularize the loss function [40, 41]. For this reason, as typically done for training neural networks, the loss function in Eq. 8 is only computed for a subset of the training data, in mini-batches of size $M$, instead of the whole training set of size $N$. At the beginning of each iteration, the subset that is used is then randomly resampled, effectively making the loss function stochastic.
Appendix B Details of the First Study
B.1 Model
To generate training data for the neural network, we combine a linear auto-regressive input, with a stochastic nonlinear output model. Specifically, we considered an input that evolves according to AR(1) statistics:
$S_{t}=\phi S_{t-1}+\xi_{t}$ (17)
where $\xi_{t}$ are iid random variables from a unit Gaussian distribution, and $\phi\in[0,1)$ is a parameter. In steady state, the autocovariance of this process is given by
$\langle S_{\tau}S_{\tau+t}\rangle=\frac{\phi^{|t|}}{1-\phi^{2}}\,.$ (18)
The output $X_{t}$ is governed by the equation
$X_{t}=\sigma(\gamma S_{t})+\rho X_{t-1}+\vartheta\eta_{t}$ (19)
@@ B.1 Model
is the logistic function. The gain $\gamma$ effectively controls the strength of the nonlinearity; see Fig. 1. This process models a response that saturates as the input grows. In fact, $\sigma(x)$ is equivalent to the Hill function commonly used in biochemistry to describe saturating enzyme kinetics [42].
B.2 Training the Neural Networks
@@
We trained our machine learning model with synthetic data generated according to Eq. 19 for various values of the gain, denoted by $\gamma$. For each value of $\gamma$, we created a distinct training set of 1000 pairs of time series $(s_{1:50},x_{1:50})$ and trained one autoregressive model per training set. Once the models were trained, we estimated the mutual information for each of them using PWS, employing variational inference to perform the marginalization.
@@
For the generative model, we use a recurrent neural network (RNN). Using Gated Recurrent Unit (GRU) cells [39] with a hidden size of 64 for temporal processing, along with a dense layer that outputs two values, representing the mean $\mu$ and log-variance $\ln\sigma^{2}$ of a Gaussian distribution. For each time step $i$, the model receives input signals $s_{i}$ and $x_{i-1}$ and predicts the next output value $x_{i}$ by sampling from this Gaussian. The model is trained by iteratively optimizing over 100 epochs over the entire training set. The training set is split into mini-batches of size 64 which are shuffled after each epoch. We optimize the model using the Adam optimizer [43] ($b_{1}=$0.9$,b_{2}=$0.999$\$) with weight decay regularization [44] of $1\text{\times}{10}^{-4}$ and using a cosine decay learning rate schedule [45] that smoothly decreases the learning rate from $1\text{\times}{10}^{-2}$ to $1\text{\times}{10}^{-3}$ throughout the training process.
The inference model used for marginalization is also modeled as a RNN. This model consists of an encoder-decoder architecture for approximating the posterior $\operatorname{\mathrm{P}}(s_{1:n}|x_{1:n})$. The encoder first processes the input sequence $x_{1:n}$ in reverse using a GRU cell with 64 hidden units to create a latent sequence $h_{1:n}$. The decoder then uses another RNN with a hidden size of 64 (using a GRU cell) to autoregressively generate a sequence $\tilde{s}_{1:n}$ using the context sequence $h_{1:n}$. The sequence $\tilde{s}_{1:n}$ is generated probabilistically by sampling from a mixture of logistic distributions with five components, as suggested in Ref. [46]. The inference network is trained for 100 epochs with mini-batches of size 64. For each $x_{1:n}$ in the training set, 16 Monte Carlo draws from the inference network $\tilde{s}_{1:n}\sim q(\tilde{s}_{1:n}|x_{1:n})$ are used to estimate the ELBO loss. The loss function gradient is estimated using SGVB [37]. The model is optimized using the ADAM optimizer with weight decay regularization (same parameters as above). We use an initial learning rate of $5\text{\times}{10}^{-3}$ that decays exponentially by a factor of $0.5$ as training progresses.
To compute the marginal probability $\operatorname{\mathrm{P}}(x_{1:n})$ from our models, we use the inference network to generate $2^{14}=$16\,384$\$ samples $\tilde{s}_{1:n}$ for each sequence $x_{1:n}$. From these samples $\tilde{s}_{1:n}$, the marginal probability is estimated using Eq. 9. We monitor the convergence of the variational marginalization procedure by computing the effective sample size from the importance weights [47]. In our case, the effective sample size always remained above 85% of the actual sample size, indicating that the inference network approximates the posterior well.
Appendix C Details of the Second Study
C.1 Dataset
@@
The stimulus is a horizontally moving vertical bar displayed on a screen. The $x$-position of the bar is varied stochastically. Its dynamics are given by a stochastically driven harmonic oscillator model. We rescale the input data, such that the variance $\langle x^{2}\rangle$ is 1. This model is defined by the following equations
$\tau\frac{dx}{dt}$ $=v$ (21)
$\tau\frac{dv}{dt}$ $=-\frac{1}{4\gamma^{2}}x-v+\frac{1}{\sqrt{2}\gamma}\eta_{v}(t)$ (22)
@@
C.2 Neuron model
In each time-interval $t_{n},t_{n+1})$ we assume the spike counts of each neuron are Poisson distributed. The firing intensity of neuron $k$ at time $t_{i}=i\Delta t$ is denoted by $\lambda^{k}_{i}$. The random variable $X^{k}_{i}$ denotes the number of spikes of neuron $k$ in the time-interval $t_{i:i+1}=[t_{i},t_{i+1})$ and is assumed to be Poisson-distributed. The probability of observing $x$ spikes in this interval is thus given by
$\mathrm{P}(X^{k}_{i}=x)=\frac{(\lambda^{k}_{i}\Delta t)^{x}}{x!}e^{-\lambda^{k}_{i}\Delta t}\,.$ (23)
Therefore, the negative log likelihood is given by
$\mathcal{L}=\sum_{i}-n_{i}\ln\lambda_{i}\Delta t+\lambda_{i}\Delta t+\ln(n_{i}!)$ (24)
@@ C.2 Neuron model
During each training step, the model receives the stimulus $s_{1:N}$ and the observed spike counts $\boldsymbol{x}_{1:N}$ at every timestep, predicts the time-varying firing intensity of each neuron, and computes the loss between the predicted and actual spike counts. The loss function used in the model is the Poisson negative log-likelihood loss. Optimization is performed using the Adam optimizer, and the learning rate is dynamically adjusted using a OneCycleLR scheduler [48], which initially increases the learning rate before gradually decaying it to improve convergence.

### v2

@@ Discussion
We demonstrated how autoregressive sequence prediction models can be trained on time-series data to learn a generative model, which can be combined with PWS to compute the information rate. By applying ML-PWS to a nonlinear model, we showed that it provides more accurate mutual information estimates than the Gaussian approximation. While this example serves as a proof of concept, it shows the potential of advanced machine learning techniques to automatically derive stochastic models from experimental data, and to enable the computation of information-theoretic measures for complex, high-dimensional data.
@@
While ML-PWS shares similarities with the DoE estimator [30], ML-PWS has the benefit that $\operatorname{\mathrm{P}}(x_{1:n})$ is obtained via an exact marginalisation of $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})\operatorname{\mathrm{P}}(s_{1:n})$ using PWS, and not by training a second neural network as in DoE; this ensures that the marginal statistics $x_{1:n}$ are consistent with the input distribution $\operatorname{\mathrm{P}}(s_{1:n})$, as determined by the setup of the experiment, and the conditional input distribution $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$ of the generative model. Moreover, ML-PWS also enables the efficient computation of the mutual information $I(\boldsymbol{S}^{\prime},\boldsymbol{X}^{\prime})$ between a input signal $\boldsymbol{S}^{\prime}$ with different statistics and the corresponding output $\boldsymbol{X}^{\prime}$, without needing to re-train the marginal model. This is particularly useful if one is interested in the channel capacity, as determined by the input distribution that maximizes the mutual information or information rate for the system of interest. In systems without feedback, $\operatorname{\mathrm{P}}(\boldsymbol{x}|\boldsymbol{s})$ is a property of the system and does not change upon changing the input. The generative model then remains the same, such that ML-PWS can directly recompute the mutual information for different input statistics.
Acknowledgments
@@
Code Availability
The code used to generate the results of this study is openly available [39].
Appendix A: Gaussian Autoregressive RNN model
@@
$h_{i}=f_{\theta}(s_{i},h_{i-1})$ (A1)
where $h_{i}\in\mathbb{R}^{d}$ for $i\in\{1,\ldots,n\}$ and an activation function $f_{\theta}:\mathbb{R}\times\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. The activation function $f_{\theta}$ could, for instance, be a fully connected neural network layer. We instead use LSTM cells [40] for $f_{\theta}$ which allow the model to better learn long-term dependencies. In any case, given an initial state $h_{0}$ and the parameters $\theta$, an RNN represents a deterministic map from one sequence $s_{1:n}$ to another sequence $h_{1:n}$.
From the sequence $h_{1:n}$ we can obtain a stochastic representation of the output sequence $x_{1:n}$. We decompose $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$ as a product of conditional probabilities
$\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})=\prod^{n}_{i=1}\operatorname{\mathrm{P}}(x_{i}|x_{1:i-1},s_{1:i})\,.$ (A2)
Specifically, we extend Eq. A1 by adding a sampling step to obtain $x_{i}$ from $s_{i}$ and $x_{i-1}$
$h_{i}$ $=f_{\theta}\left(\begin{bmatrix}s_{i}\\ x_{i-1}\end{bmatrix},h_{i-1}\right)$ (A3)
$x_{i}\mid h_{i}$ $\sim\mathcal{N}(\mu(h_{i}),\sigma(h_{i}))$ (A4)
such that each $x_{i}$ is a normally distributed random variable with mean $\mu(h_{i})$ and standard deviation $\sigma(h_{i})$. $\mu$ and $\sigma$ are typically modeled as neural networks.
The autoregressive model represents a generative model for the distribution $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$. Generating $x_{1:n}\mid s_{1:n}$ is done sequentially. For $i=1,\ldots,n$, the sampling procedure alternates between computing $h_{i}$ using Eq. A3 to get the parameters of the conditional distribution, $\hat{\mu}_{i}=\mu(h_{i})$, $\hat{\sigma}_{i}=\sigma(h_{i})$, and sampling the next $x_{i}\sim\mathcal{N}(\hat{\mu}_{i},\hat{\sigma}_{i})$. Thus, the conditional probability of the resulting output sequence is given by
$\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$ $=\prod^{n}_{i=1}\operatorname{\mathrm{P}}(x_{i}|x_{1:n-1},s_{1:n})$ (A5)
@@
In practice, we use the following form for $\mu$ and $\sigma$
$\mu(h)$ $=W_{\mu}h+b_{\mu}$ (A6)
$\sigma(h)$ $=\exp(W_{\sigma}h+b_{\sigma})$ (A7)
where $\boldsymbol{\theta}=(\theta,W_{\mu},b_{\mu},W_{\sigma},b_{\sigma})$ are model weights that are trained.
There are a few practical considerations for efficiently training the model. Training is performed in iterations and, as typically done for training neural networks, the loss function in Eq. 7 is only computed for a subset of the training data, in mini-batches of size $N_{\text{batch}}=25$, instead of the whole training set of size N=1000N=\$1000\$. At the beginning of each iteration, the subset that is used is then randomly resampled.
Appendix B: Details of the First Study
To generate training data for the neural network, we combine a linear auto-regressive input, with a stochastic nonlinear output model. Specifically, we considered an input that evolves according to $\mathrm{AR}(p)$ statistics:
$S_{t}=\sum^{p}_{j=1}\phi_{j}S_{t-j}+\xi_{t}$ (B1)
where $\xi_{t}$ are iid random variables from a unit Gaussian distribution, and the $\phi_{j}\in[0,1)$ are model parameters. The output $X_{t}$ is governed by the equation
$X_{t}=\sigma(\gamma S_{t})+\rho X_{t-1}+\vartheta\eta_{t}$ (B2)
@@
is the logistic function. The gain $\gamma$ effectively controls the strength of the nonlinearity; see Fig. 1a. This process models a response that saturates as the input grows. In fact, $\sigma(x)$ is equivalent to the Hill function commonly used in biochemistry to describe saturating enzyme kinetics [41].
Table B1: Model Parameters for Eqs. B1 and B2. We use an $\mathrm{AR}(3)$ process.
Input Output
Parameter $\phi_{1}$ $\phi_{2}$ $\phi_{3}$ $\gamma$ $\rho$ $\vartheta$
Value 0.5 -0.3 0.2 1.0 0.2 0.2
@@
For this case study, we trained our machine learning model with synthetic data generated according to Eqs. B1 and B2 for various values of the gain, denoted by $\gamma$. The other parameters are specified in Table B1. For each value of $\gamma$, we created a distinct training set of $N=1000$ pairs of time series $(s_{1:50},x_{1:50})$ and trained one autoregressive model per training set. Once the models were trained, we estimated the mutual information for each of them using PWS, employing variational inference to perform the marginalization; see also Appendix D: PWS simulation details below.
@@
For the generative model, we use a recurrent neural network (RNN). Using LSTM cells with a hidden size of 64 along with a dense layer that outputs two values, representing the mean $\mu$ and log-variance $\ln\sigma^{2}$ of a Gaussian distribution. For each time step $i$, the model receives input signals $s_{i}$ and $x_{i-1}$ and predicts the next output value $x_{i}$ by sampling from this Gaussian. The model is trained by iteratively optimizing for 100 epochs over the training set in mini-batches of size 64 which are shuffled after each epoch. We optimize the model using the Adam optimizer [42] ($b_{1}=$0.9$,b_{2}=$0.999$\$) with weight decay regularization [43] of $\lambda=$1\text{\times}{10}^{-4}$\$ and using a cosine decay learning rate schedule [44] that smoothly decreases the learning rate from $1\text{\times}{10}^{-2}$ to $1\text{\times}{10}^{-3}$ throughout the training process.
The inference model used for marginalization is also modeled as a LSTM and represents a variational approximation of the posterior $\operatorname{\mathrm{P}}(s_{1:n}|x_{1:n})$. It can be seen as an autoregressive normalizing flow that transforms a sequence of iid Gaussian random numbers $\epsilon_{1:n}$ into the samples $s_{1:n}$ [45]. At each time step $i$ it receives the current output observation $x_{i}$ and a noise sample $\epsilon_{i}\sim\mathcal{N}(0,1)$. First, $x_{i}$ is transformed by a fully connected neural network layer into a feature vector $y_{i}\in\mathbb{R}^{k}$ where we choose $k=16$. This vector $y_{i}$ is concatenated with the previous prediction $s_{i-1}$ and passed to an LSTM cell which outputs two parameters $u_{i}$ and $v_{i}$. These parameters parametrize the affine transformation $s_{i}=u_{i}\epsilon_{i}+v_{i}$. The network is trained for 200 epochs with mini-batches of size 64. For each $x_{1:n}$ in the training set, 16 Monte Carlo draws from the inference network $\tilde{s}_{1:n}\sim q(\tilde{s}_{1:n}|x_{1:n})$ are used to estimate the ELBO loss. The loss function gradient is estimated using SGVB [37]. The model is optimized using the ADAM optimizer with weight decay regularization (same parameters as above). We use an learning rate of $1\text{\times}{10}^{-2}$.
Appendix C: Details of the Second Study
@@
The stimulus is a vertically moving horizontal bar displayed on a screen. The $x$-position of the bar is varied stochastically. Its dynamics are given by a stochastically driven harmonic oscillator model. We rescale the input data, such that the variance $\langle x^{2}\rangle$ is 1. This model is defined by the following equations
$\tau\dot{x}$ $=v$ (C1)
$\tau\dot{v}$ $=-\frac{1}{4\gamma^{2}}x-v+\frac{1}{\sqrt{2}\gamma}\eta(t)$ (C2)
@@
Our ML network models neuron spike probabilities as follows. In each time-interval $t_{n},t_{n+1})$ we assume the spike counts of each neuron are Poisson distributed. The firing intensity of neuron $k$ at time $t_{i}=i\Delta t$ is denoted by $\lambda^{k}_{i}$. The random variable $X^{k}_{i}$ denotes the number of spikes of neuron $k$ in the time-interval $t_{i:i+1}=[t_{i},t_{i+1})$ and is assumed to be Poisson-distributed. The negative log likelihood is given by
$\mathcal{L}=\sum_{i}-n_{i}\ln\lambda_{i}\Delta t+\lambda_{i}\Delta t+\ln(n_{i}!)$ (C3)
@@
Optimization is performed using the Adam optimizer, and the learning rate is dynamically adjusted using a OneCycleLR scheduler [46], which initially increases the learning rate before gradually decaying it to improve convergence.
Appendix D: PWS simulation details
The central idea of ML-PWS is to develop a generative model that can learn the conditional probability distribution $\operatorname{\mathrm{P}}(x_{1:n}|s_{1:n})$ and then compute the marginal output $\operatorname{\mathrm{P}}(x_{1:n})$ via Monte Carlo averaging in trajectory space. We find that for relatively short signal trajectories, as in the application of ML-PWS to neuronal data (Fig. 2), it is sufficient to generate input trajectories $s_{1:n}$ from an importance sampling distribution $q(s_{1:n}|x_{1:n})$ using the ELBO variational procedure, and then obtain $\operatorname{\mathrm{P}}(x_{1:n})$ by directly averaging over the full trajectories using Eq. 8. In contrast, for longer signal trajectories, it is necessary to improve the sampling distribution $q(s_{1:n}|x_{1:n})$ via a sequential Monte Carlo approach, generating the trajectories and performing the concomitant marginalization segment by segment, as in RR-PWS [31]. Indeed, for the application of ML-PWS to the autoregressive model (Fig. 1), we combined RR-PWS with ELBO.